# Trust, But Verify: A Survey of Randomized Smoothing Techniques

### v1

@@ II-B Robustness-Accuracy Trade-off
A compositional architecture termed ACES, which aims to tackle this robustness-accuracy trade-off, is introduced in [49]. It decides on a per-sample basis whether to use the robust-but-less-accurate smoothed model or the accurate-but-less-robust base classifier for prediction using a learnable selection algorithm that can again be made robust by further smoothing. This approach improves the trade-off between robustness and accuracy in current models and can be used as an orthogonal optimisation to other methods. This is done in the light of optimising randomized smoothing specific robustness and accuracy using compositional architectures, as introduced and proposed earlier by Mueller, Mark Niklas et al. [93].
@@
A notable work by Mohapatra, Ko, et al. [32] articulates and proves two major limitions regarding RS: (I) the decision boundary of the smoothed classifier will shrink, resulting in a discrepancy in class-wise accuracy, and (II) applying noise augmentation in the training process does not resolve the shrinking issue because of inconsistent learning objectives. The paper proceeds to review adversarial robustness certification, exposes the significant hidden cost of RS, which includes biased predictions using evidence from both real-life and synthetic datasets, provides a comprehensive theory exposing the root of the biased prediction-shrinking phenomenon and then discusses the effects of data augmentation on the shrinking phenomenon and its implications. In essence, they arrive at two observations - the need to limit the use of high values of smoothing factor $\sigma$ and a data geometry-dependent augmentation scheme to counteract the shrinking effect caused by smoothing properly. This challenges what we already know about randomized smoothing, which, as introduced by Cohen et al., uses data-independent noise parameters. Namely, the Gaussian noise, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ which is convoluted with the base classifier is data-independent and $\sigma$ can be thought of as a hyperparameter representing a trade-off between accuracy and robustness. Too low of a value will give smaller certified radii, which are given, after appropriate approximations, as the certified radius $r=\sigma\Phi^{-1}(\underline{p_{A}})$, which is directly proportional to $\sigma$. Using too large of a value for $\sigma$ will reduce the accuracy in calculating $\underline{p_{A}}$ due to highly noisy training inputs.
@@
II-C High Inference Cost
Randomized smoothing utilizes Monte Carlo sampling, which requires multiple passes of the noisy samples through the model, leading to high inference costs, at the same time it has been shown that certified radius increases with the no. of passes or no. of samples ($n$) and for significant radius size typically $10^{5}$ or more samples are required [1]. This implies that during inference we need to do a forward pass to have a significant certified radius and essentially increases the computational cost by a large margin.
A controlled way to trade off the average certified radius (ACR) with the number of samples required was introduced by Chen, Ruoxin, et al. [23]. They changed their sampling scheme from an input-agnostic scheme to an input-specific scheme. To achieve this input-specific sampling, first a relatively loose two-sided Clopper-Pearson interval is calculated for a given input using a smaller sample size. Then, a mapping of sample size to the relative decline in average radius is computed for the input. This mapping shows a trade-off between the number of samples and the average certified radius.
Another way of reducing the computation cost could be by lowering the inference time of the smoothed model by techniques such as pruning, quantization of the base model, or knowledge transfer. But this essentially means we have changed the actual model and the theoretical guarantees may not hold. Ugare, Shubham, et al. [10] study this scenario and formulate an incremental certification technique for randomized smoothing called Incremental Randomized Smoothing (IRS). This improves performance on probabilistic techniques like randomized smoothing compared to existing deterministic techniques for incremental certification. In particular, incremental certification deals with certifying robustness bounds for the smoothed classifier of a given modification of a base classifier, reusing information from the smoothed classifier of the base classifier to reduce computational costs. It does so based on the disparity between the classifiers and certain approximation techniques. On a related note, Vaishnavi, Pratik, et al. [120] propose a similar knowledge transfer scheme, Certified Robustness Transfer (CRT), for randomized smoothing for training $l_{2}$ certifiably robust image classifiers with comparable levels of robustness using a pre-trained classifier that is certifiably robust as a teacher. Both techniques demonstrate significantly reduced costs, indicating the empirical effectiveness of their methods.
@@ II-C High Inference Cost
Challenges in the Scalability of RS In this section, we explored the challenges which affect RS. While the robustness-accuracy tradeoff, as noted, is a challenge which affects the entire field of adversarial robustness and even ML as a whole, the other two, viz., the curse of dimensionality and the issue of high inference costs, expressly limit the scalability of RS. Further, on a broad level, we note these challenges are acknowledged and understood in their impact, whether implicitly, as is true for computational costs, or explicitly, as is seen for the curse of dimensionality. However, contrary to expectations, little research is devoted to developing solutions that truly address these issues and improve the scalability of RS.
III Improving Randomized Smoothing
Randomized smoothing offers a tight lower bound on the certified radius for any classifier. As with any certified robustness technique, these bounds are lower than the upper bounds arising from empirical verification techniques, which are based on heuristics and empirical evidence instead of theoretical guarantees. Thus, there has been much research into improving the performance of randomized smoothing and improving its certified radii to an optimal limit. Further, there is a need to address the challenges faced in RS and develop various techniques. In this section, we explore specific optimisations that improve the robustness certificates given by RS and solve various issues encountered in RS.
III-A Randomized Smoothing for Training

### v2

@@ 2.2 Robustness-Accuracy Trade-off
While as noted above, there is not a dearth of works which study the theoretical robustness-accuracy trade-off, there are very few works which aim to actually address it. In this regard, we note the work on a compositional architecture termed ACES ([19]) which decides on a per-sample basis whether to use the robust-but-less-accurate smoothed model or the accurate-but-less-robust base classifier for prediction using a learnable selection algorithm that can again be made robust by further smoothing. This is done in the light of optimising randomized smoothing specific robustness and accuracy using compositional architectures, as introduced and proposed earlier in [43].
2.3 High Inference Cost
We note that high inference cost is an issue which plagues randomized smoothing. This is because it utilizes Monte Carlo sampling, which requires multiple passes of the noisy samples through the model, leading to high inference costs and for significant radius sizes, typically $10^{5}$ or more samples are required as per [8].
A controlled way to address this was to trade off the average certified radius (ACR) with the number of samples required using an input-specific scheme ([30]). Another way of reducing the computation cost could be by lowering the inference time of the smoothed model by techniques such as pruning, quantization of the base model, or knowledge transfer. But this essentially means we have changed the actual model and the theoretical guarantees may not hold. In this regard, a incremental certification technique for randomized smoothing is Incremental Randomized Smoothing (IRS) which improves performance on probabilistic techniques like randomized smoothing compared to existing deterministic techniques for incremental certification ([58]). This works by reusing information from the smoothed classifier of the base classifier to reduce computational costs. On a related note, we note Certified Robustness Transfer (CRT) as a training technique based on knowledge transfer ([59]).
While these techniques display promising empirical results, they often lack the theoretical justification behind their validity. Further, there is much still left to be explored in this regard, as can be seen by the few numbers of works written in this regard.
2.4 Challenges in the Scalability of RS
In this section, we explored the challenges which affect RS. While the robustness-accuracy tradeoff, as noted, is a challenge which affects the entire field of adversarial robustness and even ML as a whole, the other two, viz., the curse of dimensionality and the issue of high inference costs, expressly limit the scalability of RS. Further, on a broad level, we note these challenges are acknowledged and understood in their impact, whether implicitly, as is true for computational costs, or explicitly, as is seen for the curse of dimensionality. However, contrary to expectations, little research is devoted to developing solutions that truly address these issues and improve the scalability of RS.
3 Improving Randomized Smoothing
Randomized smoothing offers a tight lower bound on the certified radius for any classifier. As with any certified robustness technique, these bounds are lower than the upper bounds arising from empirical verification techniques, which are based on heuristics and empirical evidence instead of theoretical guarantees. Thus, there has been much research into improving the performance of randomized smoothing and improving its average certified radii. Further, there is a need to address the challenges faced in RS and develop various techniques. In this section, we explore specific optimisations that improve the robustness certificates given by RS and solve various issues encountered in RS.
3.1 Randomized Smoothing for Training
The idea of incorporating randomized smoothing objectives in adversarial training roots from the intuition that doing so will result in boosted performance and improved robustness. Further, addressing the challenge of training a good base classifier that is accurate and robust when smoothed using a soft classifier with a hard classifier led to improved results ([49]). Multiple works in this regard achieve larger average certified radii in less training time than state-of-the-art adversarial training algorithms ([63, 14, 24]). Further, these methods may use adaptive techniques and improve on the tightness of the robust bounds. These approaches not only are they computationally efficient and empirically sound, but also orthogonal to other empirical defense schemes. We can link these results to scalability and conclude that these methods make RS scalable).
3.2 Improving Performance of Pre-Trained Classifiers
Improving the performance of pre-trained classifiers by turning them from non-robust into provably robust ones is desired to maintain their confidentiality, integrity, and availability. This includes using computationally inexpensive denoisers as denoised smoothing can be applied to both white and black box settings, making it the state-of-the-art approach to defending pre-trained classifiers with improved methodologies ([50, 52, 29]). While the work so far in this domain seems somewhat promising, there remains scope for future research to further improve the certified accuracy of high-dimensional datasets with increasing $\sigma$.
3.3 Leveraging Local Information
The original randomized smoothing technique makes use of Monte-Carlo sampling to derive certified radii. This information is fully summarised for an input $x$ by $p_{A}(x)$, the probability of the top class, and does not require any other information beyond the standard deviation $\sigma$ of the Gaussian. There has been a series of works which explore the utilisation of the gradient information, for second-order smoothing, or higher-order derivatives to achieve tight bounds, addressing the curse of dimensionality ([31, 2, 40]).
While these works explore optimization using local information and achieve improved bounds, a drawback of this approach is that the computational overheads associated with calculating local information add heavily to the already expensive smoothing convolution with the gains they add being bound ([32]).
3.4 Ensemble-Based Methods
Randomized smoothing bounds can be improved using ensemble-based methods. These methods work by aggregating the results of multiple classifiers, somewhat similar to how RS involves smoothing the base classifier with multiple noisy inputs and using those results to create the base classifier.
Various works have shown state-of-the-art results using ensembles as the base model for RS. One insight offered in the work is that the reduced variance of ensembles over the perturbations introduced in RS leads to significantly more consistent classifiers, while still maintaining robustness ([18, 57]). Further, these papers address the inference cost issue by developing techniques to reduce the prediction and certification cost of the models, such as patch-based sampling or adaptive prediction algorithms or reducing sample complexity with data-dependent adaptive sampling ([20, 35]).
Despite the prohibitive costs of ensemble training, adaptive algorithms have been shown to reduce the computational overhead. Overall, ensemble methods provide a satisfactory approach to improve robustness guarantees with reduced costs as compared to other similarly expensive approaches.
3.5 Input-Specific Information
RS uses a data-independent hyperparameter, $\sigma$, for the Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ to be convoluted with the base classifier. Even though this forms a basis of the robustness-accuracy tradeoff, it is also important to note that the issue of an optimal choice of $\sigma$ is important to improve certified robustness using TS. To deal with this, [1] propose the use of additional information which can be extracted from the data itself which improves the choice of the hyperparameter $\sigma$ to achieve a larger certified radius. Related to this, there have been a variety of works which gradually change the structure of the covariance matrix used in RS, making it anisotropic and data-dependent and rotation-robust ([1, 7, 60, 26, 48]), albeit without theoretical guarantees and while suffering from the curse of dimensionality ([53]). Further, many techniques use information beyond the Neyman-Pearson lemma, such as classifier-specific information, data-dependent information, local geometric information and non-Gaussian distributions for noise convolution ([10, 30, 53, 3]). However, these methods only add a computational overhead, instead of working towards reducing the inference costs. In this, we point out the work on QCRS ([23]) which as a computationally efficient approach with less overhead and is one of the few input-specific methods to address the complexity issue directly.
3.6 Improving Randomized Smoothing for Scalability
Throughout our examination in this section, we observe that most new optimizations in randomized smoothing do not directly address the curse of dimensionality or the issue of high inference costs. Instead, they tend to introduce new techniques, each with its own computational overhead and dimensionality challenges, and then propose solutions for that additional cost. One such example is that of input-dependent techniques, which add input-specific components over an already costly randomized smoothing process, as pointed out in that specific subsection. While this certainly helps improve the applicability of a technique individually, it does not enhance the overall scalability of randomized smoothing, hence this foundational issue persists.
4 Scalability in Practice
When considering the practical scalability of RS, several key challenges emerge across various application domains. In high-dimensional data scenarios, such as high-resolution images or complex sensor inputs, the curse of dimensionality significantly hampers RS’s effectiveness. Although a lot of works are there that discuss this issue but as already mentioned in section 2 most of these work are theoretical in nature and only a few provide practical solutions to this issue. Even then at times those solution bring additional computational overhead.
Additionally, The high inference cost poses difficulties for real-time systems requiring rapid decision-making, like autonomous vehicles or live video analysis. In general, All the methods of RS that we have discussed are computationally very expensive, irrespective of the challenge or the improvement they are doing. Additionally, some methods put an additional computational overhead over RS as discussed in some of the sub sections of section 3. Finally comes the general issue of robustness-accuracy trade-off, which further limits the use of RS in domains which require high accuracy but at the same time high reliability such as medical domain.
4.1 Future Directions for Scalable RS
To truly make RS scalable, future research could focus on developing practical techniques with theoretical justifications focused on the following:
  1. 1.
    Efficient dimensionality handling: Developing techniques that can effectively deal with high-dimensional inputs without significant information loss or computational overhead.
  2. 2.
    Reduced inference costs: Exploring methods to dramatically reduce the number of samples required for certification without compromising on robustness guarantees.
  3. 3.
    Adaptive smoothing: Developing techniques that can dynamically adjust the level of smoothing based on the input and the required level of robustness, potentially reducing unnecessary computations.
  4. 4.
    Compression techniques: Investigating ways to compress the smoothed classifier without significant loss in robustness guarantees.
5 Conclusion
This paper has reviewed randomized smoothing (RS) as a certification-based defense against adversarial attacks in machine learning, focusing on its theoretical foundations, practical challenges, and recent advancements and scrutinizing them from the perspective of scalability. While RS offers robust theoretical guarantees, its practical application faces challenges like the curse of dimensionality and high computational costs, limiting its effectiveness in real-world, high-dimensional scenarios. Despite various improvements, such as training-based methods and ensemble techniques, these often worsen scalability issues.
The persistent robustness-accuracy trade-off requires careful consideration, with potential solutions like compositional architectures adding complexity. Future research should focus on enhancing RS’s scalability by developing efficient techniques for high-dimensional inputs, reducing sample requirements for certification, designing hardware-aware algorithms, and exploring adaptive smoothing and compression methods. In summary, while RS provides certified robustness against adversarial attacks, its scalability remains a significant barrier to widespread adoption. Addressing these challenges is essential for bridging the gap between RS’s theoretical strengths and practical deployment in large-scale applications.
Appendix A Appendix / supplemental material
A.1 Curse of Dimensionality
In their work, [8] left the case of general $l_{p}$ norms as an open problem, suspecting that smoothing with other noise distributions might provide robustness guarantees for these cases, and only discussed randomized smoothing in the context of $l_{2}$ norms. Additionally, they state that, despite the certified radius being independent of the input dimension $d$, randomized smoothing can be easily scaled to higher dimension images as in the case of higher dimension images, we can increase the variance of the noise, which will eventually lead to increase in the robust radius. This approach has its issues, as will be discussed in the next section. Later, two independent works, [4] and [22], showed that as the dimension of the input increases, it becomes increasingly difficult to defend against the general $l_{p}$ norm cases for $p>2$ as the certified radius decreases.
[4] showed that, given any noise distribution $\mathcal{D}$ and the noise vector $\epsilon\sim\mathcal{D}$, 99% pixels of the noise vector, must satisfy $\mathbb{E}\epsilon_{i}^{2}=\Omega(d^{1-\frac{2}{p}}r^{2}\frac{1-\delta}{\delta^{2}})$, where $r$ is the robust radius and $\delta$ is the difference between the probability scores of the top two classes predicted by the smoothed classifier (highest score-class $c_{A}$ and runner-up class $c_{B}$). They further proved that the problem of finding the certified radius in the $l_{\infty}$ case can be approximated as finding the $l_{2}$ radius times $1/\sqrt{d}$ of that radius.
Subsequently, [22] show, in the specific case of a generalized Gaussian distribution, tighter bounds than those attained by [4]. Their work states that the certified radius for an $l_{p}$ norm decreases as $O\left(d^{\frac{1}{2}-\frac{1}{p}}\right)$ with data dimension $d$ for $p>2$. The family of distributions discussed in this paper includes the Laplacian, Gaussian, and uniform distributions, commonly employed in literature for randomized smoothing techniques. [61] then extend the discussion for the case of spherical symmetric distributions (which includes the Gaussian distribution) and present upper bounds for $p>2$. It also discusses upper bounds for $p=2$ for many popular smoothing distributions, which have yet to be discussed. [62] go into more detail about $l_{1}$, $l_{2}$, and $l_{\infty}$ threat models to derive from the class of i.i.d. distributions, the most optimal smoothing distribution using the concept of Wulff Crystals. Overall, their results in this regard can be summarised as follows: the uniform distribution is the most optimal choice for $l_{1}$ attacks, the Gaussian distribution, as mentioned by [8], is the most optimal for $l_{2}$ attacks, and that for $l_{\infty}$, the Gaussian distribution performs best using appropriate approximations. There is a tight $O\left(\min\left(1,d^{\frac{1}{2}-\frac{1}{p}}\right)\right)$ dimensionality bound for $l_{p}$ norms without using any additional information other than the Neyman-Pearson technique.
[12] provide a different perspective and argue about the information-theoretic limitations faced by RS and how they are not intrinsic but a byproduct of the current certification methods. They further discuss that these certificates need to include more information about the classifier and ignore the local curvature of the decision boundary, which leads to sub-optimal robustness guarantees as the dimension of the problem increases. It is possible to overcome this by generalizing the Neyman-Pearson Lemma and collecting more information about the classifier. The paper shows that it is possible to approximate the optimal certificate with arbitrary precision by probing the decision boundary with several noise distributions. More specifically, the paper achieves this without sampling in high dimensions by combining uniform and Gaussian distribution and leveraging the isotropic properties of the latter. Further, this process retains natural accuracy as it is executed at certification time. The paper then gives theoretical insight into how to mitigate the computational cost of a classifier-specific certification.
[47] present notable work in the sense that they implicitly propose a solution to the curse of dimensionality, suggested by [4]. It combines the limitations of randomized smoothing in light of the curse of dimensionality with the manifold hypothesis ([13]) and shows that using various dimensionality reduction techniques improves robust guarantees in theory, as opposed to heuristic approaches in previous works by [44] and [25].
So far, only [34] have provided proofs and a possible solution to the issue of the curse of dimensionality using a technique called double sampling randomized smoothing (DSRS). They utilize a primary distribution in the form of a general Gaussian distribution and additional information in the form of a sampled probability from a secondary truncated distribution that is not necessarily different from the primary distribution. For $l_{2}$ norm, they provide a tighter radius and prove that their method can certify $\Theta\left(\sqrt{d}\right)$ robust radius bounds. As mentioned by [8], for the $l_{\infty}$ case, the certified radius is equivalent to $1/\sqrt{d}$ times the $l_{2}$ robust radius, which on using DSRS would remove the dependency on the dimensions of the input.
We note that there has been an in-depth theoretical exploration of the issue termed the curse of dimensionality. The base conclusion of all papers remains that there is a tight bound on certified radii that decreases with increasing dimensionality, which arises from statistical, probabilistic, and information-theoretic perspectives based on the existing methodology for RS. While we certainly encourage more research into that aspect, we notice the dire lack of solutions for the same. Apart from DSRS, no breakthrough has been achieved to solve this issue. As a result, we encourage exploring various solutions to this issue.
A.2 Robustness-Accuracy Trade-off
In adversarial robustness literature, not limited to RS, there is a focus on increasing the robustness bounds of a classifier. While this seems to be an optimal goal, there has been literature suggesting that increasing robustness comes with a tradeoff in accuracy ([64] and [56]). It has been studied for adversarial robustness in much detail, including theoretical considerations. There are accuracy costs of improving robustness in a model. Both empirically and theoretically, it is seen that robustness and accuracy exist in a trade-off, contrary to what one would expect intuitively. Randomized smoothing is also vulnerable to this trade-off, and there is much to explore in this regard, leaving this an important open aspect of research for randomized smoothing techniques. Intutively seen, randomized smoothing involves smoothing, the base classifier with Gaussian noise. While this provides certified robustness, it can be inferred that there is some loss in accuracy due to noisy inputs. Termed as the robustness-accuracy tradeoff, we explore this issue concerning RS in this subsection. Specifically, increasing the noise hyperparameter $\sigma$ results in a larger robustly certified radius but decreased accuracy of the model. One must note that there also exists, in general, a trade-off between accuracy and robustness in RS concerning a smaller robust radius for more accurate models, not just limited to the choice of $\sigma$.
Specific to RS, the theoretical and numerical aspects of Gaussian smoothing are explored in the work by [39], along with which they provide theoretical results showing the limitations of RS in terms of classification accuracy. Mainly, they characterize and identify the conditions under which Gaussian smoothing leads to a decrease in classification accuracy, provide theoretical lower bounds for the magnitude of this effect, numerically inspect the behaviour of the certified radius, and use tools from information theory to analyse the effects of Gaussian smoothing during training augmentation, concluding that it leads to information loss and finally validate their results empirically.
A theoretical analysis of robustness-accuracy in terms of the benign risk in accuracy, i.e., accuracy on non-adversarially-corrupted samples, associated with applying randomized smoothing on a classifier trained using noise augmentation for the base classifier is given by [45]. This arises from the observation of improved performance when using a noise-augmented base classifier instead of one without any such augmentation. They derive an upper bound for any data distribution used for smoothing, which suggests randomized smoothing harms the classifier’s accuracy. This practice has pervaded most literature with ample empirical support. The paper then suggests certain distributions where smoothing may be beneficial based on the separation regime. Also, it suggests optimisation for randomized smoothing, using different noise parameters to train the base classifier. They also discuss the theoretical intuition behind this, viz., the empirical observation that real data lies on a lower dimension than the actual data dimension.
A compositional architecture termed ACES ([19]), which aims to tackle this robustness-accuracy trade-off, decides on a per-sample basis whether to use the robust-but-less-accurate smoothed model or the accurate-but-less-robust base classifier for prediction using a learnable selection algorithm that can again be made robust by further smoothing. This approach improves the trade-off between robustness and accuracy in current models and can be used as an orthogonal optimisation to other methods. This is done in the light of optimising randomized smoothing specific robustness and accuracy using compositional architectures, as introduced and proposed earlier by [43].
@@ A.2 Robustness-Accuracy Trade-off
A notable work by [41] articulates and proves two major limitions regarding RS: (I) the decision boundary of the smoothed classifier will shrink, resulting in a discrepancy in class-wise accuracy, and (II) applying noise augmentation in the training process does not resolve the shrinking issue because of inconsistent learning objectives. The paper proceeds to review adversarial robustness certification, exposes the significant hidden cost of RS, which includes biased predictions using evidence from both real-life and synthetic datasets, provides a comprehensive theory exposing the root of the biased prediction-shrinking phenomenon and then discusses the effects of data augmentation on the shrinking phenomenon and its implications. In essence, they arrive at two observations - the need to limit the use of high values of smoothing factor $\sigma$ and a data geometry-dependent augmentation scheme to counteract the shrinking effect caused by smoothing properly. This challenges what we already know about randomized smoothing, which, as introduced by Cohen et al., uses data-independent noise parameters. Namely, the Gaussian noise, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ which is convoluted with the base classifier is data-independent and $\sigma$ can be thought of as a hyperparameter representing a trade-off between accuracy and robustness. Too low of a value will give smaller certified radii, which are given, after appropriate approximations, as the certified radius $r=\sigma\Phi^{-1}(\underline{p_{A}})$, which is directly proportional to $\sigma$. Using too large of a value for $\sigma$ will reduce the accuracy in calculating $\underline{p_{A}}$ due to highly noisy training inputs.
@@
A.3 High Inference Cost
Randomized smoothing utilizes Monte Carlo sampling, which requires multiple passes of the noisy samples through the model, leading to high inference costs, at the same time it has been shown that certified radius increases with the no. of passes or no. of samples ($n$) and for significant radius size typically $10^{5}$ or more samples are required ([8]). This implies that during inference we need to do a forward pass to have a significant certified radius and essentially increases the computational cost by a large margin.
A controlled way to trade off the average certified radius (ACR) with the number of samples required was introduced by [30]. They changed their sampling scheme from an input-agnostic scheme to an input-specific scheme. To achieve this input-specific sampling, first a relatively loose two-sided Clopper-Pearson interval is calculated for a given input using a smaller sample size. Then, a mapping of sample size to the relative decline in average radius is computed for the input. This mapping shows a trade-off between the number of samples and the average certified radius.
Another way of reducing the computation cost could be by lowering the inference time of the smoothed model by techniques such as pruning, quantization of the base model, or knowledge transfer. But this essentially means we have changed the actual model and the theoretical guarantees may not hold. [58] study this scenario and formulate an incremental certification technique for randomized smoothing called Incremental Randomized Smoothing (IRS). This improves performance on probabilistic techniques like randomized smoothing compared to existing deterministic techniques for incremental certification. In particular, incremental certification deals with certifying robustness bounds for the smoothed classifier of a given modification of a base classifier, reusing information from the smoothed classifier of the base classifier to reduce computational costs. It does so based on the disparity between the classifiers and certain approximation techniques. On a related note, [59] propose a similar knowledge transfer scheme, Certified Robustness Transfer (CRT), for randomized smoothing for training $l_{2}$ certifiably robust image classifiers with comparable levels of robustness using a pre-trained classifier that is certifiably robust as a teacher. Both techniques demonstrate significantly reduced costs, indicating the empirical effectiveness of their methods.
@@ A.3 High Inference Cost
A.4 Randomized Smoothing for Training